### v2

# How Much Do LLMs Hallucinate across Languages? On Multilingual Estimation of LLM Hallucination in the Wild

Abstract
In the age of misinformation, hallucination—the tendency of Large Language Models (LLMs) to generate non-factual or unfaithful responses—represents the main risk for their global utility. Despite LLMs becoming increasingly multilingual, the vast majority of research on detecting and quantifying LLM hallucination are (a) English-centric and (b) focus on machine translation (MT) and summarization, tasks that are less common “in the wild” than open information seeking. In contrast, we aim to quantify the extent of LLM hallucination across languages in knowledge-intensive long-form question answering. To this end, we train a multilingual hallucination detection model and conduct a large-scale study across 30 languages and 6 open-source LLM families. We start from an English hallucination detection dataset and rely on MT to generate (noisy) training data in other languages. We also manually annotate gold data for five high-resource languages; we then demonstrate, for these languages, that the estimates of hallucination rates are similar between silver (LLM-generated) and gold test sets, validating the use of silver data for estimating hallucination rates for other languages. For the final rates estimation, we build a knowledge-intensive QA dataset for 30 languages with LLM-generated prompts and Wikipedia articles as references. We find that, while LLMs generate longer responses with more hallucinated tokens for higher-resource languages, there is no correlation between length-normalized hallucination rates of languages and their digital representation. Further, we find that smaller LLMs exhibit larger hallucination rates than larger models.
How Much Do LLMs Hallucinate across Languages?
On Multilingual Estimation of LLM Hallucination in the Wild
Saad Obaid ul Islam1 Anne Lauscher2 Goran Glavaš1 1WüNLP, CAIDAS, University of Würzburg {saad.obaid-ul-islam,goran.glavas}@uni-wuerzburg.de 2Data Science Group, University of Hamburg anne.lauscher@uni-hamburg.de
1 Introduction
Generalizing seamlessly to (seemingly) arbitrary language understanding, reasoning, and generation tasks, Large Language Models (LLMs) (Kojima et al., 2022; Dubey et al., 2024; Aryabumi et al., 2024; Yang et al., 2024) have arguably become the first ubiquitously adopted language technology, with application ranging from search engines (Xiong et al., 2024), interactive agents (Teubner et al., 2023) and knowledge retrieval (Yu et al., 2023) to various content generation tasks (Liu et al., 2022b). Their utility, however, is hindered by their tendency to hallucinate (Maynez et al., 2020; Zhou et al., 2021; Ji et al., 2023; Zhang et al., 2023), that is, produce information that is either (i) inaccurate or factually incorrect with respect to the objective state of the world (e.g., in open-ended question answering) or (ii) unfaithful with respect to some reference (e.g., in summarization).
Consequently, a large body of work on tackling LLM hallucination has emerged, with efforts falling into the three main areas: (1) detection, i.e., identification of the hallucinated content; (2) evaluation, primarily focusing on measures for quantifying the extent and severity of hallucinations ; and (3) mitigation, focusing on mitigating hallucinative tendencies of LLMs (Ji et al., 2023). While significant progress has been made in English (Maynez et al., 2020; Liu et al., 2022a; Obaid ul Islam et al., 2023; Chen et al., 2023; Lin et al., 2022; Kasai et al., 2024; Mishra et al., 2024; Vu et al., 2023), hallucination evaluation efforts targeting other languages have been much sparser Clark et al. (2023); Guerreiro et al. (2023); Herrlein et al. (2024); Shafayat et al. (2024). Moreover, these efforts have primarily targeted high-resource languages (Qiu et al., 2023; Shafayat et al., 2024) with benchmarks limited to reference-based tasks—text summarization (Clark et al., 2023; Aharoni et al., 2022) and machine translation (Dale et al., 2023; Guerreiro et al., 2023). While highly relevant, these tasks are arguably less representative of LLM usage ‘in the wild’ Trippas et al. (2024), where asking knowledge-intensive questions (and expecting free-form answers) is more prominent.
In this work, we address the above gaps in multilingual hallucination detection and evaluation research with the ultimate goal of estimating the “in the wild” hallucination rates of LLMs across languages. Multilingual estimation of such hallucination rates is challenging due to the scarcity of multilingual hallucination benchmarks covering open-ended knowledge-seeking tasks that are representative of real-world LLM usage: unlike in reference-based generation tasks like summarization and machine translation, LLMs often generate long-form responses to open-ended questions, requiring more comprehensive evaluation approaches (Wei et al., 2024b). Concretely, we present a large-scale study that estimates hallucination rates for 30 languages (both high(er)- and low(er)-resource languages). Our main contributions are as follows: (1) We train a multilingual hallucination detection model; to this end, we resort to the translate-train approach Artetxe et al. (2023); Ebing and Glavaš (2024), i.e., we machine-translate an English hallucination-labeled training dataset to all other languages; (2) We create (i.e., gold) hallucination evaluation datasets with span-level fine-grained human annotations for five high-resource languages and generate synthetic (i.e., silver) hallucination evaluation datasets for 25 additional languages; (3) We propose a protocol for estimating “in the wild” hallucination rates of LLMs in a massively multilingual setting; (4) We report hallucination rate estimates for six open-source LLM familes and 30 languages, finding that hallucination rates decrease with model size and increase with model language support. To the best of our knowledge, this work represents the first attempt to estimate hallucination rates of LLMs in the wild: (i) across many languages and on (ii) knowledge-intensive question answering, representative of real-world LLM usage. Figure 1 provides an overview of our whole framework111We release our datasets and work on: https://github.com/WorldHellow/mHallucinations-LLM.
2 Background and Related Work
We provide a brief overview of the body of related work on (1) hallucination detection models and (2) benchmarks for evaluating LLM hallucination.
Hallucination Detection. Coarsely, LLM hallucinations fall into two categories. Intrinsic hallucination are content contradicts some reference information source. The reference may be explicitly given to the LLM as part of the task (e.g., the text to be summarized in summarization or source language text in machine translation) or it may implicit (e.g., general world knowledge in question answering). In contrast, extrinsic hallucination refers to content that does not contradict the reference but is unnecessary or superfluous with respect to the task (e.g., additional facts in fact-based question answering) (Ji et al., 2023). Recent work introduced finer-grained taxonomies for both categories. For example, Mishra et al. (2024) distinguish between several types of intrinsic hallucinations (e.g., entity-based hallucinations or relation-based hallucinations). In a similar vein, extrinsic hallucinations are split into subtypes such as invented, subjective, and unverifiable content.
Unsurprisingly, most hallucination detection (and classifications) models are based on neural languages models. These are are either pre-trained encoder LM (Zhou et al., 2021; Liu et al., 2022b), discriminatively fine-tuned to classify texts as containing hallucinations or not or LLMs prompted (zero-shot or with in-context examples) to detect hallucinations (Manakul et al., 2023; Yang et al., 2023) or fine-tuned to generate hallucinated spans (Mishra et al., 2024). In this work, we cast hallucination detection as a span-detection task, formulated discriminatively, with a classifier on top of an “encoder-based” LM. However, instead of resorting to small pretrained encoder LMs, we bidirectionally (i.e., discriminatively) fine-tune a larger generative LLM, following recent advances in converting decoder LMs into encoders Li et al. (2023b); Dukić and Šnajder (2024); BehnamGhader et al. (2024); Schmidt et al. (2024).
Hallucination Benchmarks. Hallucination detection models as well as evaluation datasets have largely focused on English vary in the granularity from document-level (Yang et al., 2023) of annotations/predictions, over passage- and sentence-level annotations (Zhou et al., 2021; Manakul et al., 2023), to fine-grained token- or span-level annotations (Liu et al., 2022a; Mishra et al., 2024). Notable examples include SelfCheckGPT (Manakul et al., 2023), HaluEval (Li et al., 2023a), and ScreenEval (Lattimer et al., 2023), which measure hallucination detection rates in summarization and single-fact question answering. Multilingual benchmarks for evaluating hallucination detection models remain sparse and focus on reference-based tasks like machine translation (Dale et al., 2023) and summarization (Qiu et al., 2023) which poorly represent the LLM usage in the wild.
Faithfulness in reference-based tasks is complemented by truthfulness (i.e., factuality) in question answering. Most benchmarks, e.g., TruthFulQA Lin et al. (2022), RealtimeQA Kasai et al. (2024), FreshQA Vu et al. (2023), and SimpleQA Wei et al. (2024a) here are English-centric and cover only questions that require a simple single-factoid answer. LongFact (Wei et al., 2024b), Factscore (Min et al., 2023) and mFactScore Kim et al. (2024) do test LLMs truthfulness in generating long and free-form answers. However, LongFact is an English-only benchmark, whereas Factscore and mFactscore, albeit multilingual, cover a very specific domain of biographic questions.
3 Hallucination Detection
We first describe how we obtained multilingual hallucination detection datasets (§3.1) and then report on training and evaluation of a multilingual hallucination detection model (§3.2).
3.1 mFAVA Benchmark
Evaluation Datasets. We start from the English FAVA (Mishra et al., 2024) dataset and its respective set of fine-grained hallucination types. FAVA’s evaluation portions were created by (1) eliciting information-seeking prompts (i.e., questions) from various sources,222See the original paper for more details. (2) generating responses with three LLMs and (3) having human annotators label hallucinated spans. We follow a similar protocol to create evaluation datasets for 30 languages.333§A.1 Figure 7 lists the mFAVA languages. We start from 300 information-seeking prompts, 150 from evaluation portion of FAVA and 150 from the Natural Questions dataset (Kwiatkowski et al., 2019). We then ask GPT-4 Achiam et al. (2023) to (1) first create answer passages in a target language and then to (2) explicitly introduce the hallucinations of the fine-grained FAVA types into the answer. We refer to these synthetically labeled hallucination evaluation datasets, comparable across the 30 target languages, as mFAVA-Silver.
For five linguistically diverse high-resource languages—Arabic, Chinese, German, Russian, and Turkish—we also collect human hallucination annotations. To this end, we provide to the annotators the reference Wikipedia page, and the (hallucination-enriched) generation from mFAVA-Silver (of course, without the GPT-4’s hallucination annotations). We source the annotations via Prolific, recruiting 5 annotators per language: all five annotators first annotated the same 50 instances, after which they were given non-overlapping sets of 50 more instances. We provide more details on the annotation process (and costs) in the §A.2. We measure the inter-annotator agreement (IAA) in terms of pairwise-averaged Cohen’s kappa on token-level class decisions, both with (IAA Category) and without (IAA Binary) considering the fine-grained hallucination types. As shown in Figure 2, we observe satisfactory to good IAA for all five languages. Regarding the 50 instances labeled by all annotators, we ultimately take the annotations of the annotator that has the highest IAA with hallucination annotatios of GPT-4 from mFAVA-Silver. We denote the final human-labeled evaluation datasets for the five high-resource languages with mFAVA-Gold. Figure 2 also shows the overall IAA between human annotations from mFAVA-Gold and GPT-4’s synthetic annotations from mFAVA-Silver (Silver-Gold): interestingly, we observe that human annotators on average agree more with GPT-4 than with one another.
| Very Unlikely | Unlikely | Neutral | Likely | Very Likely |
| 21.8% | 24.7% | 13.0% | 25.3% | 15.2% |
Because we synthesize the hallucinated content with GPT-4 (the annotators, of course, did not know that nor which part of the generation was meant to be a hallucination according to GPT-4), there is a risk that these hallucinations may not be realistic in the sense that they can fool a human reader. Because of this, we asked our annotators to additionally indicate (on a 5-degree Likert scale from “very unlikely” to “very likely”) the likelihood of hallucination fooling a human reader for each span that they labeled. Table 1 reveals that more than half of the labeled hallucinations were judged as convincing (i.e., not unlikely to fool a human). The silver test set statistics for all 30 languages are shown in §A Figure 6. Gold annotations statistics are shown in Table 2.
Training Dataset. The FAVA training set, consisting of ca. 30K instances, is fully synthetically created in the same way as the test portion, just without the human annotation step. We automatically translate the training portion of the FAVA dataset using NLLB (Costa-jussà et al., 2022) to our 30 target languages. After translation, we project the span-level annotations to token-level labels using the simple Inside-Out (I-O) scheme (Ramshaw and Marcus, 1995)444In preliminary experiments, we also tested the B-I-O scheme, but I-O led to better span detection performance.. Like our evaluation benchmark mFava, we prepare training data for two tasks: (1) detecting hallucinated spans, regardless of hallucination type (Binary task: tokens are classified as either part of a hallucinated span or not) and (2) detection and hallucination type classification (Category task: 7-way classification, tokens classified into one of 6 FAVA hallucination types or as not part of a hallucinated span).
| ENT | REL | INV | CON | UNV | SUB | Total | |
| RU | 184 | 65 | 188 | 287 | 211 | 153 | 1,088 |
| AR | 144 | 10 | 171 | 123 | 150 | 69 | 667 |
| ZH | 264 | 18 | 259 | 282 | 265 | 139 | 1,227 |
| DE | 546 | 25 | 311 | 324 | 333 | 238 | 1,777 |
| TR | 149 | 27 | 288 | 244 | 161 | 149 | 1,018 |
| Total | 1,287 | 145 | 1,217 | 1,260 | 1,120 | 748 | 5,777 |
| German | Chinese | Arabic | Russian | Turkish | ||||||||
| Task | Model | Context | Silver | Gold | Silver | Gold | Silver | Gold | Silver | Gold | Silver | Gold |
| Mono | Bidirect | 78.0 | 58.0 | 62.4 | 55.1 | 75.3 | 54.4 | 78.9 | 60.7 | 78.5 | 66.7 | |
| Binary | Multi | Bidirect | 89.5 | 65.0 | 69.7 | 58.7 | 82.5 | 61.6 | 89.1 | 65.5 | 86.4 | 72.5 |
[…]
| Mono | Bidirect | 53.4 | 38.3 | 35.2 | 22.6 | 14.6 | 7.3 | 63.3 | 36.2 | 49.1 | 30.3 | |
| Category | Multi | Bidirect | 73.2 | 45.0 | 46.5 | 30.1 | 66.1 | 37.2 | 72.3 | 41.5 | 72.9 | 51.8 |
[…]
3.2 Multilingual Hallucination Detection
Models. Using the translations of ca. 30K FAVA training instances in our 30 target languages, we train the following models: (1) Mono denotes monolingual models trained on data of one language (and evaluated for the same language on the respective mFAVA portion), i.e., we train 30 Mono models, one for each of our target languages; (2) Multi refers to a single multilingual model trained on concatenated training data of all 30 languages. We train the Mono models and the Multi model for both tasks, Binary and Category. We follow the recent body of work that successfully converts generative decoder LLMs into encoders for discriminative tasks Li et al. (2023b); Dukić and Šnajder (2024); BehnamGhader et al. (2024); Schmidt et al. (2024) and fine-tune Llama-3-8B-base Dubey et al. (2024) by removing future-token masking, i.e., allowing for bidirectional contextualization (Bidirect). For comparison, for the Multi model, we also fine-tune the decoder as-is, using the default causal token masking (i.e., unidirectional contextualization; Causal).
Training. In all cases, we freeze the original model parameters and train QLora adapters (Dettmers et al., 2024), with three runs (random seeds) for each experiment, reporting mean performance. The input to the models is the reference Wikipedia article, prepended to the LLM-generated answer to the user question/prompt, with the cross-entropy loss computed exclusively over the tokens of the LLM-generated answer. We provide further training details in §A.3 for training details.
Results. Table 3.1 summarizes the hallucination detection performance for five high-resource languages for which we have both LLM-synthesized Silver data and human-annotated Gold portions in our mFAVA benchmark. We first observe that, expectedly, just detecting hallucinated spans (Binary task) is much easier than additionally correctly recognizing the type of hallucination (Category task). Although category labels offer finer-grained insight into the nature of LLM hallucination, we deem the models’ performance on fine-grained hallucination type classification—especially on Gold, human-labeled portions of mFAVA—insufficient for reliably estimating type-specific hallucination rates “in the wild” (see §4.2). These results are in line with IAA from Figure 2, with consistently larger IAA for hallucination detection (Binary) then for type classification (Category). This renders fine-grained hallucination type classification difficult for both humans and models and warrants a broader research effort on hallucination type taxonomies as well as better hallucination type detection models. We leave this for future work.
Models’ performance on the detection-only (Binary) tasks is much better across the board, but the results are much better on the Silver portions (hallucinations generated by GPT-4) of mFAVA than on the Gold (human-labeled hallucination spans). This is expected, because the hallucinated spans in our training data have also been generated by GPT-4—this means that the human-annotated Gold mFAVA portions introduce much more of a distribution shift w.r.t. training data than the corresponding Silver portions. At this point it is important to (re-)emphasize that we are not really interested in the absolute performance of the detection models, but rather using these detection performance estimates to produce reliable hallucination rate estimates for LLMs in the wild (§4.2).
We next observe that the 30-language multilingual model (Multi) is consistently better than language-specific monolingual models (Mono), with gaps being particularly wide in the Category task (e.g., +30 F1 points for Arabic on the Gold mFAVA portion). Albeit smaller, the differences are also substantial in the Binary hallucination detection (e.g., +7 F1 points for Arabic and German, on respective Gold mFAVA portions). Finally, bidirectional contextualization in fine-tuning (Bidirect) seems to be generally more effective than fine-tuning with future-token masking (Causal), with Chinese performance as the only exception. This is in line with findings from other token-classification tasks Li et al. (2023b); Dukić and Šnajder (2024).
4 Estimating Hallucination in the Wild
We next propose a protocol for estimating hallucination rates of LLMs (for a wide range of languages) in the wild, based (1) on the number of hallucinated tokens detected by a hallucination detection (HD) model in the wild and (2) estimates of HD model’s performance (precision and recall).
4.1 From Model Performance to Hallucination Rates Estimates
Estimating Hallucination Rates in the Wild. Let and be the estimates of token-level precision and recall of a HD model for some language and let be the number of hallucination tokens that the HD model detected (i.e., predicted) on some corpus of LLM generations in language , which serves as an approximation of the LLM outputs in the wild. We then posit that the estimate of the true hallucination rate of the LLM in the wild for language , , is given as follows:
where is the total number of tokens in , i.e., the total number of tokens generated by the LLM across answers to all user prompts. Intuitively, multiplying the number of model’s detections with its estimated precision discounts by the number of tokens falsely detected as hallucinated by the model—while we do not know exactly which token predictions are false positives, the expected rate of false positives is, by definition, exactly captured by . Analogously, dividing with accounts for the tokens that are hallucinated, but will (falsely) not be detected by the model—and is exactly the estimate of the rate of such false negatives. We divide the estimate of the absolute number of truly hallucinated tokens (i.e., ) with , making a relative measure, that is, a rate (i.e., proportion) of all generated tokens that are hallucinated (multiplied by 100 and expressed as %). We provide a more detailed explanation/justification of Eq. (1) in §A.4.
Estimation Dataset. We next create corpora (one corpus for each of our 30 target languages) of free-text LLM answers to knowledge-intensive queries, as approximations of the LLM usage in the wild. We start by randomly selecting articles from the language-specific Wikipedia, to serve as ground truth reference text. To ensure quality of reference text, we choose only from Wikipedia articles that are at least 2,000 characters long and have the collaborative Wikipedia depth (Alshahrani et al., 2023) of at least 5.555The depth indicates the number of collaborative edits and correlates with the quality/factuality of the content. We then prompt GPT-4 to generate two knowledge-intensive queries for each selected article, ensuring that the information required to answer to the query is fully contained in the article text (see Table 6 in the §A.5 for the exact prompt). As a sanity check, we manually checked for 50 synthesized queries and five languages from Table 3.1—by translating the query and reference article to English—whether the answers to queries are indeed contained in the article, establishing that this is indeed so in 98% of cases. Our final dataset for multilingual hallucination rate estimation consists of 25,685 Wikipedia articles (spanning over 15,940 unique Wikipedia categories) and 51,133 queries. Table 9 in §A provides per-language statistics. We provide details on constructing the datasets in §A.5.
Finally, we collected responses to all queries from a total of 11 instruction-tuned open-source LLMs from 6 families (ranging in parameter count from 2 to 9 billion): Llama-3.x (Dubey et al., 2024), Aya-23 (Aryabumi et al., 2024), Euro-LLM (Martins et al., 2024), Gemma-2 (Team, 2024) Qwen-2.5 (Yang et al., 2024), and Mistral v3 (Jiang et al., 2023). We divided the queries into five subsets: for each subset the LLMs generated responses with a different random seed (see Table 8 for details on the generation configurations).
Estimates from mFAVA-Silver Performance. On the one hand, creating Gold datasets for hallucination detection evaluation is prohibitively expensive (see §A.2)—this is why we obtained such annotations for only five of 30 mFAVA languages. On the other hand, the estimates of HD model’s performance are much higher on mFAVA-Silver (see Table 3.1), with GPT-4-labeled hallucinations: this, at first glance, questions the validity of estimating ‘in the wild’ hallucination rates based on and estimated on Silver data, for the 25 languages for which we do not have mFAVA-Gold portions. Recall, however, that we do not care about HD model’s absolute and , but whether the and estimates can produce reliable hallucination rate estimates . Looking at Eq. 1, depends on the ratio and not absolute values of and . We thus next test, for the five languages with both Silver and Gold portions in mFAVA, whether the estimates based on the Silver and (roughly) match those based on Gold and . Figure 3 shows estimates, computed from the performance of our Multi (Bidirect) model on Silver and Gold portions, respectively, and number of its hallucination detections on outputs of three LLMs: Llama-3-8B, Qwen-2.5-7B, and Aya-8B. We observe very strong Pearson correlation between the Gold-based and Silver-based estimates, which, we argue, justifies the usage of Silver mFAVA datasets for estimating for the 25 languages without the Gold mFAVA portions.
4.2 Final Estimates
Figure 4 shows our in-the-wild hallucination rate estimates for all 30 mFAVA languages and 11 LLMs. The average rate across all languages varies between 7% and 12%, with both Gemma models offering the lowest rates. Smaller Qwen-2.5 (3B) model hallucinates the most and, interestingly, significantly more than its larger counterpart (9B).
More Parameters, Less Hallucination? For each LLM, we have 15 estimates (3 HD model instances 5 generations by the LLM) of (averaged across all languages): we apply the Student’s t-test to determine if the differences between models (smaller and larger) significantly differ. Figure 5(a) summarizes the results. The difference between the two EuroLLM variants is not significant; larger Gemma model hallucinates more (significantly), but the HR are low for both variants; for Llama and Qwen, the smaller models hallucinate significantly more. Finally, we aggregate the estimates across all “small” models (1.7-3B) and all “large” models (7-9B) and see that, overall (column “Overall” in Figure 5(a)), smaller LLMs hallucinate significantly more (). This agrees with Wei et al. (2024b) who report larger models to be more truthful in long-form answer generation.
More Languages, More Hallucination? Figure 5(b) compares LLMs’ hallucination rates against their declared number of supported languages. We see that LLMs that support more languages tend to hallucinate more (e.g., EuroLLM supports 35 languages, whereas Gemma is declared to support English only)—the correlation is strong and significant ().
Say Less, Hallucinate Less? Intuitively, one would expect LLMs’ hallucination rates to be larger for languages in which they are less competent (i.e., seen the least in pretraining and instruction-tuning). Surprisingly, however, we do not find this to be the case for any of the models. E.g., we observe the lowest hallucination rate for Sindhi ( of tokens are hallucinated), a language with merely 18,000 Wikipedia articles and largest hallucination rate for Hebrew (). Across all 30 languages, however, we find no correlation between the hallucination rates and measures of language “resourceness”: (i) proportion of language-specific data in Common Crawl and (ii) number of articles in the language-specific Wikipedia. As illustrated in Figure 5(c), we do observe that LLMs generate longer responses for languages in which they are more competent—this entails a larger number of hallucinated tokens for longer responses, but not (necessarily) a larger (per-token) hallucination rate (recall that we account for the response length in Eq. 1). Indeed, we observe no correlation whatsoever between the response length and hallucination rates across languages (). This suggests that a trade-off between the answer length and the amount (not rate!) of hallucinations is a largely language-independent property of LLMs.
5 Conclusion
We presented the first effort towards understanding how much multilingual LLMs hallucinate “in the wild”. To this end, we proposed a novel framework for hallucination rate estimation, which adjusts the number of detected hallucinations based on the detector’s performance resulting in more reliable rate estimates. We trained a series of multilingual detection models, and measured their precision and recall scores on our newly created mFAVA datasets across 30 languages. To estimate hallucinations, we build a novel synthetic open-domain knowledge-intensive QA dataset for which we collected answers from eleven open-source LLMs. Our findings indicate that smaller models and models that cover more languages hallucinate significantly more, and that model response-length does not correlate with hallucination rate.
Limitations
We acknowledge the following limitations of our work: (1) we adopted the common translate-train approach and thus used MT to translate the original FAVA to our 30 target languages. While one may argue that we thus add some noise to the training process resulting in unreliable detectors, recall that we are not opting for the highest possible detection performances, but rather interested in obtaining reliable performance estimates. (2) We only have gold annotations for 5 languages. Here, one might argue that, thus, our performance estimates might be unreliable. This is why in §4.1, we compare estimates obtained on mFAVA-Silver with ones obtained on mFAVA-Gold and show that silver annotations can serve as a reliable proxy. (3) For our hallucination evaluation, we only manually check a subset of the Arabic, Chinese, German, Russian, and Turkish queries to ensure that the answers to the synthetic prompts are present in the Wikipedia references. The high rate of 98% we observed makes us confident that the potential error we introduce via such “non-grounded” questions for other languages is negligible. We still acknowledge, however, that the Wikipedia articles we use might be limited in terms of the knowledge they cover (Kim et al., 2024), this is why we carefully filter via minimum length and collaborative Wikipedia depth towards higher-quality articles with high coverage. (5) Finally, we deliberately limited the scope of this work to assessing factual correctness and we do not cover factual coverage. We decided to do so as quantifying hallucinations in long-form generation is already difficult for English Xu et al. (2023); Min et al. (2023); Wei et al. (2024b) and more so in non-english languages (Kim et al., 2024), and currently, resources for assessing multilingual factual coverage are still lacking.
Acknowledgments
This work was supported by the Alcatel-Lucent Stiftung and Deutsches Stiftungszentrum through the grant “Equitably Fair and Trustworthy Language Technology” (EQUIFAIR, Grant Nr. T0067/43110/23). The authors gratefully acknowledge the computing time granted by the John von Neumann Institute for Computing (NIC) and provided on the supercomputer JURECA (Jülich Supercomputing Centre, 2021) at Jülich Supercomputing Centre (JSC).
Appendix A Appendix
A.1 Choice of languages
Initially, we wanted to cover all 14 language families based on Glottolog 5.0 (Nordhoff and Hammarström, 2012)), however, as we progressed through the languages, we found that even the best closed-source LLMs like GPT-4 and Gemini are bad at generating text in low-resource languages (e.g. Amharic, Aymara, Hausa and Tamil) and we could not employ LLMs to generate and annotate a silver hallucination detection dataset in these languages. See Table 7 for 30 languages.
A.2 Annotation Process
Cost of Silver Annotations
The total cost for generating silver data for 30 languages using GPT-4 was $2,310 with $77 per language. Distribution of categories across 30 languages is provided in Table 4 and and per language label distribution is provided in Figure 6.
| ENT | REL | INV | CON | UNV | SUB | |
| Count | 11143 | 9036 | 5649 | 4024 | 5670 | 6396 |
Gold Annotations:
The annotators were sourced through prolific platform. Each annotator was screened on 10 samples and if they met the threshold of 40% agreement with the silver annotation, they were invited to participate in the full study.
It is worth noting that as the hallucination annotation task for longform QA is very cognitively demanding, it took us a long time to find annotators who could do the task correctly with high-effort. Most of the time, annotators who passed the screening test, decided to leave the study because of the high effort requirement of the task even though our study was paying above minimum wage (14 $/hr) for the full study.
The total cost of the gold annotations was (including platform and annotation fees) $4581 where each annotator was paid 14 $/hr. All the annotators were at least bachelor’s level and bilingual because in addition to understanding their own language (e.g. Arabic) they also needed to understand the task instructions and Wikipedia content (in English). The task instructions are given in Figure 8. Each annotator was asked if they consent to storing their prolific IDs during manual and automatic assessment stage. Following the assessment, their prolific IDs were deleted.
We will release mFava data under an open scientific licensing.
A.3 Training Details
All the classifiers were trained utilizing the Bi-LLM (Li et al., 2023b) and transformers (Wolf, 2019) library. The models were trained with three seeds (42, 47, 49) on 4xH100 until convergence. Seeds are set for torch.manual_seed() and random.seed(). The exact hyper-parameters are given in the Table 5. Total GPU hours: 1134.
| Parameter | Value |
| Translate Train-Val Split | 70:30 |
| Seeds | [42, 47, 49] |
| Quantization | 4-bit BF16 |
| Model | Llama-3-8B (base) |
| GPUs | 4 H100 |
| LoRA | 32 |
| LoRA | 32 |
| LoRA Dropout | 0.05 |
| LoRA Target Modules | All |
| Epochs | 2 (until convergence) |
| Input Length | 4096 |
| Learning Rate | |
| Weight Decay | 0.01 |
| Batch Size | 8 |
| Gradient Accumulation | 8 |
A.4 Adjusting for and
The hallucination rate for a given language , is defined as the ratio of hallucinated tokens detected by the model () to the total number of generated tokens ():
To refine this rate, we adjust for the detection model’s precision () and recall (). Precision is defined as:
where () denote true and false positives respectively. Rearranging this equation gives the number of true positives:
Recall is defined as:
where denotes false negatives. The total number of corrected hallucinations () can thus be expressed as:
A.5 Hallucination Evaluation Dataset
To construct the hallucination evaluation dataset, we aimed to scrap 1000 articles per language with more than 2000 characters. However, problem with non-English languages (especially moderate-low resource) is that character articles can be scarce. Furthermore, sometimes Wikipedia has articles tagged as Unreferenced, Failed Verification, or Under Construction which flag the article as unfinished or not factually verified. Such tags are very prominent in languages other than English and we do not include such articles in our dataset.
We use Wikipedia article summary (text before the first heading) as references and prompt gpt-4 to generate 2 knowledge-intensive queries per article. Sometimes, it generated only one query even though we explicitly state to generate two queries. We did not prompt the GPT-4 again to generate the second query due to budget constraints. The total cost to generate prompts for 31 languages is $192. Per language statistics can be found in Table 9. We will release hallucination evaluation data under an open scientific licensing.
Given the following reference in language <language name>: Generate two knowledge-intensive queries in <language name>. Ensure the questions are concise but knowledge-intensive. The questions should require thorough reading of the reference text to answer. Separate the questions with a newline.
A.6 Response Collection for Hallucination Evaluation Dataset
We collect LLM responses on 5 seeds: 42, 43, 44, 47, 49 for 6 LLM model666We comply with licensing agreement for each of the LLMs we use.. The generation configurations that we used are provided in the generation_config.json in model repositories on huggingface. Seeds are set for torch.manual_seed() and random.seed().
A.7 Manual Analysis
To further check the quality and informativeness (See §A.8 for definitions of informativeness) of the responses, we manually analyze 60 responses from Aya-23-8B for German and Arabic, two languages for which we have gold annotations. Overall, 10% of the responses had repetitive words and sentences and 5% of the responses were I don’t know responses. For the remainder of the samples, the responses were fluent and long and were relevant to the input prompt.
A.8 Informativeness
Currently, there is no agreed-upon definition of informativeness. Lin et al. (2022) considers a response to be informative if it is potentially relevant to the question and Wei et al. (2024b) considers a response to be informative if it has a certain number of supporting facts from the reference text.
| Language | Language Family | Script | Test-Set |
| Arabic | Afro-Asiatic (Semitic) | Arabic | Gold |
| Chinese | Sino-Tibetan (Sinitic) | Chinese (Han) | Gold |
| German | Indo-European (Germanic) | Latin | Gold |
| Russian | Indo-European (Slavic) | Cyrillic | Gold |
| Turkish | Turkic (Common Turkic) | Latin | Gold |
| Basque | Language Isolate | Latin | Silver |
| Cantonese | Sino-Tibetan (Sinitic) | Chinese (Han) | Silver |
| Catalan | Indo-European (Romance) | Latin | Silver |
| Czech | Indo-European (Slavic) | Latin | Silver |
| Esperanto | Constructed | Latin | Silver |
| Finnish | Uralic (Finnic) | Latin | Silver |
| French | Indo-European (Romance) | Latin | Silver |
| Hebrew | Afro-Asiatic (Semitic) | Hebrew | Silver |
| Hindi | Indo-Aryan | Devanagari | Silver |
| Hungarian | Uralic (Ugric) | Latin | Silver |
| Indonesian | Austronesian (Malayo-Polynesian) | Latin | Silver |
| Italian | Indo-European (Romance) | Latin | Silver |
| Japanese | Japonic | Kanji | Silver |
| Korean | Koreanic | Hangul | Silver |
| Latin | Indo-European (Italic) | Latin | Silver |
| Lithuanian | Indo-European (Slavic) | Latin | Silver |
| Malay | Austronesian (Malayo-Polynesian) | Latin | Silver |
| Polish | Indo-European (Slavic) | Latin | Silver |
| Portuguese | Indo-European (Romance) | Latin | Silver |
| Romanian | Indo-European (Romance) | Latin | Silver |
| Serbian | Indo-European (Slavic) | Cyrillic | Silver |
| Sindhi | Indo-Aryan | Arabic | Silver |
| Spanish | Indo-European (Romance) | Latin | Silver |
| Urdu | Indo-Aryan | Arabic | Silver |
| Vietnamese | Austroasiatic (Vietic) | Latin | Silver |
| Model | max_new_tokens | temperature | top_p | top_k | repetition_penalty | do_sample |
| Llama-3.x | 1024 | 0.6 | 0.9 | – | – | True |
| Aya | 1024 | – | 0.3 | – | – | True |
| Qwen-2.5 | 1024 | 0.7 | 0.9 | 20 | 1.05 | True |
| Mistral | 1024 | – | – | 50 | – | True |
| Gemma-2 | 1024 | – | – | – | – | True |
| EuroLLM | 1024 | – | – | – | – | True |
| Language | Unique Categories | Total Articles | Total Queries |
| Arabic | 537 | 959 | 1907 |
| Basque | 486 | 938 | 1872 |
| Cantonese | 261 | 401 | 793 |
| Catalan | 359 | 989 | 1976 |
| Chinese | 712 | 977 | 1939 |
| Czech | 720 | 988 | 1975 |
| Esperanto | 608 | 956 | 1912 |
| French | 332 | 987 | 1973 |
| Finnish | 549 | 995 | 1972 |
| German | 797 | 984 | 1967 |
| Hebrew | 660 | 999 | 1991 |
| Hindi | 153 | 186 | 367 |
| Hungarian | 745 | 992 | 1964 |
| Indonesian | 457 | 958 | 1913 |
| Italian | 678 | 988 | 1974 |
| Japanese | 667 | 999 | 1991 |
| Korean | 539 | 747 | 1488 |
| Latin | 334 | 465 | 916 |
| Lithuanian | 711 | 946 | 1888 |
| Malay | 442 | 778 | 1556 |
| Polish | 889 | 1000 | 1998 |
| Portuguese | 390 | 955 | 1909 |
| Romanian | 351 | 811 | 1618 |
| Russian | 462 | 999 | 1996 |
| Spanish | 938 | 977 | 1952 |
| Serbian | 386 | 798 | 1587 |
| Sindhi | 224 | 519 | 1029 |
| Turkish | 660 | 856 | 1650 |
| Urdu | 567 | 878 | 1749 |
| Vietnamese | 326 | 660 | 1311 |
| Total | 15,940 | 25,685 | 51,133 |
| Language | Precision (%) | Recall (%) | F1 Score (%) |
| GOLD | |||
| Arabic (Gold) | 73.98 | 53.40 | 61.63 |
| Chinese (Gold) | 70.73 | 53.93 | 58.79 |
| German (Gold) | 58.19 | 74.06 | 65.05 |
| Turkish (Gold) | 79.67 | 66.95 | 72.57 |
| Russian (Gold) | 63.18 | 68.46 | 65.53 |
| Average | 69.15 | 63.36 | 64.71 |
| SILVER | |||
| Arabic | 93.28 | 74.81 | 82.59 |
| Chinese | 80.33 | 66.28 | 69.77 |
| German | 91.64 | 87.77 | 89.50 |
| Turkish | 89.58 | 83.92 | 86.43 |
| Russian | 93.05 | 86.04 | 89.15 |
| Basque | 87.22 | 74.46 | 79.80 |
| Cantonese | 78.49 | 49.40 | 56.12 |
| Catalan | 94.70 | 87.46 | 90.85 |
| Czech | 93.99 | 84.75 | 89.00 |
| Esperanto | 94.28 | 86.53 | 90.05 |
| French | 91.58 | 89.37 | 90.31 |
| Finnish | 86.67 | 84.26 | 85.15 |
| Hebrew | 82.75 | 32.97 | 44.19 |
| Hindi | 68.01 | 68.48 | 66.77 |
| Hungarian | 92.35 | 74.29 | 81.93 |
| Indonesian | 92.12 | 85.75 | 88.72 |
| Italian | 93.76 | 87.26 | 90.28 |
| Korean | 86.39 | 79.11 | 82.31 |
| Japanese | 77.06 | 61.03 | 67.15 |
| Lithuanian | 90.48 | 75.39 | 81.81 |
| Malay | 86.15 | 68.96 | 75.73 |
| Portuguese | 95.80 | 86.77 | 90.94 |
| Serbian | 86.16 | 76.75 | 79.91 |
| Sindhi | 82.00 | 69.38 | 74.36 |
| Spanish | 95.86 | 85.34 | 90.14 |
| Vietnamese | 89.35 | 84.57 | 86.71 |
| Urdu | 88.82 | 72.32 | 79.39 |
| Average | 88.22 | 76.42 | 80.71 |